OpenStudy (anonymous):

Can someone help me with this one please?(: Using the replacement set, find the solutions for the equation. y = 0.5x + 10 Replacement Set: {(-2, 9.5), (0, 10),(3, 11.5),(4, 13)}

OpenStudy (anonymous):

you would just plug in x for each and if it doesnt equal the y its wrong, for example: y=0.5x+10, plug in the first pain y=0.5(2)+10 y=1+11 y=11, which is not true. and you would just do that for the rest.

OpenStudy (anonymous):

Thank you @Hope_nicole